1. Fe2O3 + 3CO → 2Fe + 3CO2 How many grams of iron metal can be refined from the mineral hematite (iron oxide) if 59.5 mol of carbon monoxide was allowed to react with it. 
1
Expert's answer
2021-03-11T08:45:27-0500

"m_{Fe}=59.5mol_{CO}\times\frac{2mol_{Fe}}{3mol_{CO}}\times\frac{55.85g_{Fe}}{1mol_{Fe}}=2.22\cdot10^3g"


Answer: "2.22\cdot10^3g"